OpenStudy (anonymous):

Simplify: (4a+3b)-(-2a-3b)

OpenStudy (anonymous):

4a+2a+3b+3b = 6(a+b)

OpenStudy (anonymous):

6a+6b=6(a+b)

OpenStudy (anonymous):

is there any exponents in there?

OpenStudy (anonymous):

no